OpenStudy (anonymous):

Find the vertices, foci and eccentricity of the ellipse. 9x^2 + y^2 = 81

OpenStudy (anonymous):

Rewrite it in standard form.

OpenStudy (anonymous):

You don't need me when you have polpak :)

OpenStudy (anonymous):

\[\frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{b^2} = 1\]

OpenStudy (anonymous):

You should divide both sides by 81, you will have: \[{x^2 \over 9}+{y^2 \over 81}=1\]

OpenStudy (anonymous):

Then rewrite 9 and 81 as the square of a and b, then you know the lengths of the major and minor axes.